A university bookstore recently sold a wirebound graph-paper notebook for $2.50, and a college-ruled notebook for $2.30. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $118.60. How many of each type were sold?

I am no good with word problems. I got the answer of 25 and 27 but could not give the equation and my teacher said it was wrong. I don't know.

Answer by scott8148(6628)

"a combination of 50 of these notebooks" ___ g + c = 50 ___ g = 50 - c

"were sold for a total of $118.60" ___ (2.50)g + (2.30)c = 118.60

substitute and solve